Operations Report – December 3rd

Crew 200 Operations Report 03-Dec-2018

SOL: 2

Name of person filing report: Antoine Bocquier

Non-nominal systems: generator autostart
Notes on non-nominal systems: nothing to report

Generator (hours run): From 4:30 pm last night to 3:30 pm this afternoon (due to snow and mud, impossible to go on EVA or engineering tour). On tonight at 5:50 pm.

Solar— SOC 90 %
Diesel Reading – N/A %
Propane Reading – N/A %
Ethanol Free Gasoline – N/A gallons

Water (auxillary tank) – 0 gallons
Water (static tank) – 500 gallons
Auxillary to Static tank transfer – no
Gallons transferred: 0
Water in GreenHab – not noted gallons, getting low
Water (loft) – Static to Loft Pump used – yes
Water Meter: not recorded

Toilet tank emptied: no

Deimos rover used: no, still not functional
Hours:
Beginning charge:
Ending charge:
Currently charging: no

Sojourner rover used: No
Hours:
Beginning charge:
Ending charge:
Currently charging: yes

Spirit rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging: yes

Opportunity rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging: yes

Curiosity rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging: yes

Notes on rovers: Rovers were not used or moved since yesterday, in charge from yesterday afternoon.
# Hours the ATVs were Used today: 0
Notes on ATVs: ATVs were taken out of active service and are now parked away from the station

HabCar used and why, where? No
CrewCar used and why, where? No

General notes and comments: Due to strong snow and mud conditions, EVA and engineering tour were not allowed. This was approved by Shannon.
Thus some of the measurements could not be performed (and generator stayed on during day, we will prevent this happening next days).

Summary of internet: After issues yesterday to connect to the network, we rebooted the Modem and everything is fine.

Summary of suits and radios: Ventilation system of suits 7 and 10 are not working. I will have a more detailed look at them tomorrow.

Summary of Hab operations: media visit today with recordings inside the hab and in the other parts of the station.

Summary of GreenHab operations: some plants planted.

Summary of ScienceDome operations: 2 microscopes may not be working, will be investigated

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Musk Solar observatory – Nominal. Robotic Observatory – Carried out maintenance: 1. Installed
replacement UPS battery, error message on startup “Warning State – Replace Battery”. Restored connections to original config, UPS switched off / plugged in. 2. Checked C14 for condensation (none significant observed) applied gentle heat to scope body for 1 hour & unplugged Tempest fan.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report [end]